find the intercept
1. y=x^3-27
2. y=x^2-3x-4

Answer by nerdybill(7384)

To find the x-intercept:
set y=0 and solve for x.
.
To find the y-intercept:
set x=0 and solve for y.
.
1. y=x^3-27
To find the x-intercept:
set y=0 and solve for x.
y=x^3-27
0=x^3-27
27=x^3
3=x
x-intercept is at (3,0)
.
To find the y-intercept:
set x=0 and solve for y.
y=x^3-27
y=(0)^3-27
y=0-27
y=-27
y-intercept is at (0,-27)
.
2. y=x^2-3x-4
To find the x-intercept:
set y=0 and solve for x.
y=x^2-3x-4
0=x^2-3x-4
0=(x-4)(x+1)
x-intercepts at (4,0) and (-1,0)
.
To find the y-intercept:
set x=0 and solve for y.
y=x^2-3x-4
y=(0)^2-3(0)-4
y=0-0-4
y=-4
y-intercept is at (0,-4)